OpenStudy (anonymous):

hey guys can I have some help please? :) What is the pH of a 1 ´ 10-5 M KOH solution? a. 3 b. 5 c. 9 d. 11

OpenStudy (anonymous):

Log(1x10^-5) = pOH = 5pH = 14 - pOH = 14 - 5 = 9 KOH is basic, and that your answer should be >7...